hartnn (hartnn):

Each letter represents a unique number from 0 to 9. Find the numbers corresponding to letters W,R,O,N,G such that it satisfies the relation given below: WRONG +WRONG -------- RIGHT

OpenStudy (lgbasallote):

my mom always told me two wrongs don't make a right....

OpenStudy (lgbasallote):

i guess i should go rob that thief who mugged me last night.

hartnn (hartnn):

well, in mathematics , multiply two negatives, u will get positive.

OpenStudy (lgbasallote):

LEFT + LEFT ==== RIGHT

hartnn (hartnn):

that's simpler,i guess,we get T directly coz T+T=T implies T=0

OpenStudy (ghazi):

WRONG+WRONG= RIGHT ....let's say W=2, R=5, O=0 N=0, G=0

OpenStudy (ghazi):

2500+2500=5000

hartnn (hartnn):

the letters are unique.

OpenStudy (ghazi):

okay...wait

OpenStudy (lgbasallote):

i suppose it's safe to assume 2(WRONG) = RIGHT

hartnn (hartnn):

^^ will that make it simpler?

OpenStudy (lgbasallote):

20000W + 2000R + 200O + 20N + 2G = RIGHT

OpenStudy (anonymous):

haha!! @lgbasallote using sixth sense...

OpenStudy (lgbasallote):

20000W + 2000R + 200O + 20N + 2G = 10000R + 1000I + 100G + 10H + T

hartnn (hartnn):

@lgbasallote may not be the best way

hartnn (hartnn):

i have 2G=T,so t is even

OpenStudy (chihiroasleaf):

w=3, r = 7, O = 3, N = 4, G = 6 37346 37346 _______ + 74692

hartnn (hartnn):

W<5

hartnn (hartnn):

@chihiroasleaf nice try,but u getting I=N

OpenStudy (lgbasallote):

then let N = 3 lol

OpenStudy (chihiroasleaf):

ahhh..., yes..., I did a mistake..., I just realize that I found I = N :(

hartnn (hartnn):

and W=O

OpenStudy (anonymous):

w=2 , R=4 , O=7 , N=6 , G=5 R=4, I=9, G=5 , H=3 AND T=0

hartnn (hartnn):

where did u get the solution??

OpenStudy (anonymous):

I just did it ....

OpenStudy (chihiroasleaf):

trial and error?

hartnn (hartnn):

show work?

OpenStudy (anonymous):

Ya... its a trial and error and a bit of logic

hartnn (hartnn):

what logic?

OpenStudy (experimentx):

W <= 4

OpenStudy (anonymous):

Ya w<=4

OpenStudy (anonymous):

And R=2W

hartnn (hartnn):

R can be 2W+1, carry over 1 can be there.

OpenStudy (anonymous):

Nope, it cannot be

hartnn (hartnn):

if R>4,then there will be carry

OpenStudy (experimentx):

WRONG WRONG ------- R I G HT -------- 2G = T <-- T should be 0, 2, 4, 6, 8

hartnn (hartnn):

thats what i already written,W<5 and T even

hartnn (hartnn):

@sauravshakya why R can't be 2W+1 ??

OpenStudy (anonymous):

Another possibility could be: W = 4, R = 9, O = 1, N = 5, G = 3 R = 9, I = 8, G = 3, H = 0, T = 6. Since carrying the added value is inevitable, there are probably more solutions also.

hartnn (hartnn):

how did u get that? trial and error or some specific way?

OpenStudy (anonymous):

Trial and error, mostly. Along with the described method above, W = 1/2 R, etc.

hartnn (hartnn):

so i will assume there is no other specific way to solve such thing, okkk thanks for the reply though.